OpenStudy (anonymous):

A) Find the average rate of change of the area of a circle with respect to its radius r as r changes from (i) 5 to 6 \rightarrow Average rate of change = (ii) 5 to 5.4 \rightarrow Average rate of change = (iii) 5 to 5.1 \rightarrow Average rate of change = (B) Find the instantaneous rate of change when r = 5 . Instantaneous rate of change =

OpenStudy (anonymous):

from five to six \[\pi \times 6^2-\pi \times 5^2=36\pi -25\pi = 11\pi\]

OpenStudy (anonymous):

the rest are similar

OpenStudy (anonymous):

what about the last part on how to find instantaneous rate of change when r = 5

OpenStudy (freckles):

Instantaneous means you need to find the derivative of the area of a circle w.r.t r (since it is the r that is changing and you are to find it at r=5)

OpenStudy (freckles):

A=pi*r^2